Circular No. 5802

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PKS 0528+134 The Compton Observatory EGRET Team (P. L. Nolan, J. Chiang,

J. Fierro, Y. C. Lin, and P. F. Michelson, Stanford University; D. L. Bertsch, B. L. Dingus, C. E. Fichtel, J. A. Esposito, R. C. Hartman, S. D. Hunter, J. R. Mattox, P. Sreekumar, and D. J. Thompson, Goddard Space Flight Center, NASA; D. A. Kniffen, Hampden-Sydney College; E. Schneid, Grumman Aerospace Corporation; and K. Brazier, G. Kanbach, H. A. Mayer-Hasselwander, and C. von Montigny, Max-Planck-Institut für Extraterrestrische Physik) communicates: “The current EGRET observation of the galactic anti-center region, which started on May 13 and ends May 25, indicates no detection of high-energy (E > 100 MeV) γ-ray emission from the quasar PKS 0528+134. Based on the first 4 days of data, the estimated 84-percent confidence upper limit is at least 2.5 times below the observed flux during Mar. 23-29 (IAUC 5753), which provided the most intense observation of this source. Together with the observations during the early part of the GRO mission (1991 May 16-30; Hunter et al. 1993, Ap.J. 409, 134), the current measurement provides strong evidence for time variability of γ-ray emission from this object. Simultaneous observations at other wavelengths would be extremely interesting.”

NOVA AQUILAE 1993 M. Mikołajewski and B. Wikierski, N. Copernicus University, Toruń,

communicate: “We estimate that a maximum brightness of V = 7.66 occurred on May 19.66 UT. During our 2-hr monitoring on May 20/21, the nova seems to show periodic variability in UBV with a period of ∼ 40-50 min and an amplitude of 0.10-0.15 mag. The following photometric data are averaged over several tens of minutes: May 19.00, V = 7.76, U–B = –0.13, B–V = +0.80, V–R = +0.96, V–I = +1.74; May 19.04, 7.71, –0.14, +0.78, +0.96, +1.73; May 20.03, 7.69, –0.01, +0.83, +0.88, +1.51; May 21.01, 8.01, –0.18, +0.64, +1.00, +1.72.”

V992 SCORPII (NOVA SCORPII 1992) Visual magnitude estimates by P. F. Williams, Heathcote, N.S.W.:

1992 Aug. 11.50 UT, 8.4; 23.44, 8.8; 31.46, 9.8; Sept. 13.44, 10.5; 1993 May 11.51, 12.8; 19.48, 12.4; 20.47, 12.6.

1993 May 22 Daniel W. E. Green